What is the perimeter of a rectangle with one side that's measurement is 45.5 and the other side is 38.5
Answer by stanbon(75887)

What is the perimeter of a rectangle with one side that's measurement is 45.5 and the other side is 38.5
--------
Perimeter = 2(L + W)
P = 2(45.5+38.5)
-----
P = 2(84)
----
Perimeter = 168 units
===========================
Cheers,
Stan H.
==============